# Rethinking Cross-Modal Fine-Tuning: Optimizing the Interaction between Feature Alignment and Target Fitting

### v1

@@ 5 Related Works
Due to limited space, we provide a short review of most relevant work on in-modal and cross-modal fine-tuning. A broader review is deferred to Appendix C.
@@ Appendix H Hyperparameters & Implementation Details
To ensure fair comparisons, we adopt the same hyperparameters as those used in ORCA (Shen et al., 2023) for model fine-tuning. These specific parameter settings are shown in Table 8 for Nas-Bench-360 (Tu et al., 2022) and Table 9 for PDEBench (Takamoto et al., 2022). For detailed implementation, we construct the target model $M_{\tau}\triangleq(\phi,p_{\tau}(z^{\prime}\mid\phi(\boldsymbol{x}^{\prime})))$ by adapting a pre-trained foundation model $M_{s}\triangleq(\theta,p_{s}(z\mid\theta(\boldsymbol{x})))$ which is selected to be a Swin Transformer (Liu et al., 2021b) for 2D tasks; and a RoBERTa (Liu et al., 2019) for 1D tasks. To compute the FLD for a regression task, we discretize the continuous label space into 10 distinct classes, consistent with the hyperparameter settings used in ORCA (Shen et al., 2023). All experiments are run on an NVIDIA RTX 2080 GPU and results are averaged over 5 independent runs. Our code repository can be found at https://github.com/khiembk/RECRAFT.

### v2

@@ 4.4 Impact of Minimizing Semantic Gap
4.5 Systematic Analysis of Existing Methods
The measurable terms FA and FLD in the theoretical bound on the target generalized loss in Theorem 7 provide a principled diagnostic tool for understanding the strengths and failure modes of prior cross-modal adaptation techniques, which remain largely empirical. In this section, we provide additional experiments to measure the FA (Eq. 14), FLD (Eq. 7), and TF (Eq. 6) terms incurred by previous cross-modal fine-tuning methods.
Figure 4: Comparison of FA, FLA, and FA across different cross-modal fine-tuning methods on the Cosmic dataset.
Based on the results illustrated in Fig. 4, we can draw the following conclusions: NFT (Naive Fine-Tuning) incurs much larger FA and FLD losses than the other methods and consequently achieves the worst performance (see Tab. 2 and Tab. 1). ORCA effectively reduces FA but barely reduces FLD by an insignificant amount. This is not surprising given that ORCA performs FA and TF in separate stages without accounting for their incurred FLD. This can make TF overcompensate for a suboptimal FA, which increases the risk of overfitting. Consequently, its effectiveness remains limited compared to our approach in most experiments. PARE has a better balance between FA and FLD compared to ORCA, but its significant reduction on FLD appears to come at a cost of increasing FA (compared to ORCA) which leads to an overall marginal reduction in FA + FLD. The significant reduction on FLD can be attributed to PARE’s design, which aims to minimize a linear combination of source and target losses using an intermediate distributional representation that combines the most important information from both the source and target modalities. Including most important information from the source modalities in the aligned representation helps constrain the feature alignment (FA) within the most important regions in the source’s representation landscape which in turn limits how far the aligned features can deviate from the source’s feature-label structure. This consequently helps avoid large FLD. However, this scheme was not optimized to prevent negative transfer, where some source-specific important information is irrelevant to the target’s task but might still be included in the intermediate representation (See Fig. 2). As such, the feature alignment must account for such extra, irrelevant information and hence, increases the FA loss. This explains why PARE incurs larger FA loss than ORCA despite having a much smaller FLD. Our method RECRAFT implements a principled minimization procedure for both FA + FLD (Section 3.1) and TF (Section 3.2), which leads to a significant reduction in both FA and FLD while also reducing TF, thus achieving better performance than all the above baselines across most experiments.
@@ 5 Related Works
Due to limited space, we provide a short review of the most relevant work on in-modal and cross-modal fine-tuning. A broader review is deferred to Appendix C.
@@ 5.2 Cross-Modal Fine-Tuning of FMs
6 Limitations and Future Works
In future work, our proposed method will be further generalized to address two existing limitations. These limitations do not undermine the soundness and practical significance of our work but addressing them can further improve its overall effectiveness. First, while the use of Euclidean cost metric to characterize the geometric structure under which the Wasserstein distance is computed does not invalidate the bound in Theorem 7, it is unclear whether a Wasserstein distance parameterized with Euclidean cost metric would optimally tighten the upper-bound. We will investigate this aspect via developing a metric learning component which parameterizes and learns this cost metric as additional parameters of the upper-bound. Second, while upper-bounding FLD with the conditional entropic terms in Eq. 18 and approximating them with a well-established pseudo-label approach (Nguyen et al., 2020) results in a valid empirical upper-bound, it is unclear whether the bound gap can be made vanishingly small as we increase the no. of sampled pseudo labels (see Eq. 19). This remains an open question to be addressed in future generalization of the current work.
@@ 7 Conclusion
Acknowledgement
The authors would like to thank the anonymous reviewers for their insightful suggestions and constructive comments, which helped to improve the quality and clarity of this paper significantly.
@@ Appendix H Hyperparameters & Implementation Details
To ensure fair comparisons, we adopt the same hyperparameters as those used in ORCA (Shen et al., 2023) for model fine-tuning. These specific parameter settings are shown in Table 8 for Nas-Bench-360 (Tu et al., 2022) and Table 9 for PDEBench (Takamoto et al., 2022). For detailed implementation, we construct the target model $M_{\tau}\triangleq(\phi,p_{\tau}(z^{\prime}\mid\phi(\boldsymbol{x}^{\prime})))$ by adapting a pre-trained foundation model $M_{s}\triangleq(\theta,p_{s}(z\mid\theta(\boldsymbol{x})))$ which is selected to be a Swin Transformer (Liu et al., 2021b) for 2D tasks; and a RoBERTa (Liu et al., 2019) for 1D tasks. To compute the FLD for a regression task, we discretize the continuous label space into 10 distinct classes, consistent with the hyperparameter settings used in ORCA (Shen et al., 2023). All experiments are run on an NVIDIA RTX 2080 GPU and results are averaged over 5 independent runs. Our code repository can be found at https://github.com/khiembk/RECRAFT